# HMT: Hierarchical Memory Transformer for Long Context Language Processing

### v2

@@ 3.3 Hierarchical Memorization
Human memory can be categorized into three strata: sensory memory, short-term memory, and long-term memory (Burgin, 2011). Sensory memory refers to very short-term memory generated from sensory information, such as vision and hearing. Short-term and long-term memory are long-lasting memories, differentiated by how long they persist in the brain. HMT is inspired by this memory hierarchy.
@@
Short-term Memory: HMT will encode the segment into an embedding that serves as a “summarization” of the segment. First, HMT will append and prepend memorization prompt embedding $H^{S}_{n}$ to the augmented segment, guiding the backbone model to encode the segment and relevant information into a summarization embedding. As depicted in Step 4 of Figure 1, we train HMT such that
$\text{BBM}(H^{S}_{n}\circ H_{n-1}[L-k,L)\circ H_{n}\circ H^{S}_{n})=D\circ H^{out}_{n}\circ H^{mem}_{n}$
@@
where $H_{n}^{mem}$ is the memory embedding of the $n^{th}$ segment. $H^{out}_{n}$ is a collection of $L$ hidden embeddings that will be used to generate logits.
@@
Long-term Memory: Each generated memory embedding will be cached as the long-term memory. Based on the memory constraint of the hardware, HMT stores the most recent N memory embeddings. The cached embeddings will be utilized as the input of memory recall mechanism to generate the memorization token embedding $H^{S}_{n}$ for each segment as illustrated in the previous section.
Note that the computation of short-term memory has some similarity to RMT (Bulatov et al., 2022), but HMT will employ the memorization prompt embeddings generated in the memory recall to augment the segments.
4 Training and Fine-tuning HMT
Since HMT introduces parameters for memory recall, we need to train new parameters and fine-tune the parameters of the backbone model cooperatively. Training HMT involves multiple segments of tokens to learn how to encode input tokens and retrieve information properly. This section will demonstrate the efficiency of HMT in long-context training and the multi-stage strategy to increase the training speed.
4.1 Long-context Training
Both HMT and RMT are trained using backward propagation through time (BPTT) (Mozer, 2013), a technique utilized to train the RNN model by unrolling recurrent forward passes of the model to optimize long-sequence learning. One issue with RMT training with BPTT is the gradient explosion and vanishing problem. With a higher BPTT unroll depth, the effectiveness of RMT will first increase and then decrease, with a slow reduction or even an increase in training loss. Figure 2 depicts the effectiveness of RMT with the OPT 2.7B backbone model assessed over the Wikitext-103 dataset with various BPTT unroll depths. For both 2k and 10k token inputs, we observe a rising PPL when unrolling more than 5 segments during training.
On the contrary, HMT will not suffer from gradient vanishing or explosion as BPTT unroll depth increases due to the memory recall mechanism. Table 2 reveals that HMT can improve its effectiveness continuously as the BPTT unroll depth increases during training. Therefore, HMT will be more effective when the BPTT unroll depth increases. A detailed gradient stability analysis is presented in Appendix H. Furthermore, we applied several techniques to optimize the GPU memory consumption to increase the maximum trainable BPTT unroll depth compared with RMT, described in Appendix I.
Table 2: Relationship between the BPTT unroll depth and the test PPL of Wikitext-103 for OPT 2.7B with HMT. The experiment is evaluated on samples with 10k tokens. HMT preserved 32 tokens from the previous segment as the sensory memory and saved 300 memory embeddings for the memory recall. The segment size is 256 tokens.
BPTT Unroll Depth Test PPL (Wikitext)
2 9.36
5 9.15
15 8.20
Figure 2: Effectiveness of training RMT with BPTT with different unroll depths for 2000 tokens and 10000 tokens input from the Wikitext-103 dataset. The backbone model is OPT 2.7B, with 256 tokens per segment during inference.
4.2 Multi-stage Training
Training HMT involves two stages. In the first stage, The model is trained without the memory recall mechanism employing BPTT with 2 segments unrolled. BPTT saves the model checkpoint locally. Then, the memory recall mechanism loads and extends the train model in the second stage. At this point, BPTT trains the modified model by unrolling the maximum number of segments that the GPUs can handle to maximize the effectiveness, which is 15 in our experiment. Since the architecture of HMT is complex, breaking up training into two stages is beneficial for local optimization and improves long context inference performance compared with single-stage training. Figure 3 exhibits the performance difference between the multi-stage and single-stage training of OPT 2.7B model with HMT for long-context inputs. Since Stage 1 involves a shorter training sequence length and a simpler recurrent architecture than Stage 2, training with Stage 1 is faster per iteration (1.15 s/iteration) than Stage 2 (3.36 s/iteration). Within the same number of training steps, multi-stage training obtains better effectiveness and lower total training time than single-stage training.
Figure 3: Training HMT + OPT 2.7B with the memory recall mechanism in two steps results in a better performance than using the mechanism to train HMT directly: first train the model without the memory recall, then load the checkpoint and train with the memory recall mechanism. Total training time is 902 s for multi-stage training and 1680 s for single-stage training on 4 AMD MI210 GPUs.

### v3

@@ 3.3 Hierarchical Memorization
Human memory can be categorized into three strata: sensory memory, short-term memory, and long-term memory Burgin (2011). Sensory memory refers to very short-term memory generated from sensory information, such as vision and hearing. Short-term and long-term memory are long-lasting memories, differentiated by how long they persist in the brain. HMT is inspired by this memory hierarchy.
@@
Short-term Memory: HMT will encode the segment into an embedding that serves as a “summarization" of the segment. First, HMT will append and prepend the memorization prompt embedding $P_{n}$ to the augmented segment. This guides the backbone model to compress the segment and relevant context into a summarization embedding with awareness of the relative positions of contexts. As depicted in Step 4 of Figure 1, we train HMT such that
$H=\text{BBM}(P_{n}||H_{n-1}[L-k,L)||H_{n}||P_{n})$ (4)
$H^{out}_{n}||M_{n}=H[k+1,L+k+2)$ (5)
@@
where $M_{n}$ is the memory embedding of the $n^{th}$ segment. $H^{out}_{n}$ is a collection of $L$ hidden embeddings that will be used to generate logits.
@@
Long-term Memory: Each generated memory embedding will be cached as the long-term memory. The cached embeddings will be utilized as the input of the memory retrieval mechanism to generate the memorization token embedding $P_{n}$ for each segment as illustrated in the previous sections.
@@ 4 Experiment
We benchmark HMT with a variety of backbone models including SmolLM 135M Allal et al. (2024), OPT 350M, OPT 2.7B Zhang et al. (2022), OpenLlamaV2 3B Geng and Liu (2023), RWKV 3B Peng et al. (2023a), and Llama 2 7B Touvron et al. (2023), under the same memory constraint (i.e. same maximum context window). Moreover, we test several models targeting long contexts (Mamba 370M Gu and Dao (2023), Yi-6B-200K Young et al. (2024), and Mistral 7B Jiang et al. (2023)) to demonstrate the benefit HMT has on generation quality and memory consumption. We evaluate HMT with state-space models (RWKV and Mamba) as backbones since we believe that models which can already process infinitely long inputs would benefit even further from HMT. All models mentioned are trained and assessed on 4 AMD MI210 GPUs, which can handle models up to 7B parameters. We further test HMT on 4 NVIDIA A100-80GB GPUs for the Qwen 2.5 14B model Bai et al. (2023a) to justify its scalability to larger models and gain a consistent effectiveness boost. To tune the extra parameters introduced by HMT, we use the RedPajamaV2 Computer (2023) dataset to pre-train each model. Notice that HMT introduced new model hyperparameters on top of the backbone model ($L$, $j$, $N$, and $k$). A common configuration is $L=1024$, $j=512$, $N=300$, and $k=32$, and we adjust these values for each model to achieve the best performance. To compare with previous works (RMT, LongMem, Memorizing Transformer, CCM), we apply the same backbone models if the method is applicable to any model, or find a backbone model with a similar size if the method requires special architecture.
For the long-context benchmark, we select subsets (NarrativeQA, Qasper, and MultiFieldQA-en for single document QA; HotpotQA, 2WikiMQA, and MuSiQue for multi-document QA; GovReport, QMSum, and Multi-News for summarization; TriviaQA for few-shot learning) from a widely acknowledged benchmark, LongBench Bai et al. (2023b), and measure them against models reported in the LongBench leaderboard. However, the maximum average document length of test sets in LongBench is shorter than 20k words, which is not very long for modern long-context models. To better understand HMT’s long-context processing ability under various context scenarios, we further study HMT on crafted and controllable dataset samples. For crafted datasets, we derive from existing datasets to form long inputs. For general language tasks, models are tested for next token generation tasks with Wikitext-103 Merity et al. (2016) (2-3k words per sample) and PG-19 Rae et al. (2019) datasets (69k words per sample on average). Samples will be concatenated or split into chunks to form longer samples and investigate the relationships between input length and the effectiveness of the model. For question-answering tasks, we chose PubMedQA Jin et al. (2019), which is a biomedical question-answering dataset with corresponding contexts. We artifact the dataset to assess HMT with multi-context inputs, described in Appendix I.
Figure 2: Test Perplexity of HMT, RMT, and three baseline models (OPT 2.7B, RWKV 3B, OpenLlamaV2 3B) with the Wikitext-103 dataset. HMT outperforms RMT by 13.0% for OPT and 10.8% for OpenLlamaV2. For RWKV, HMT can even boost the effectiveness by 16.5%, while RMT worsens the effectiveness.
Figure 3: Test Perplexity of HMT, RMT, and three baseline models (OPT 2.7B, RWKV 3B, OpenLlamaV2 3B), evaluated over the PG-19 dataset. HMT outperforms RMT by 3.98% for OPT and 6.85% for OpenLlamaV2. For RWKV, HMT can improve the effectiveness by 9.96%.
5 Results and Key Observations
In this section, we illustrate the main result of HMT. More ablation studies are in Appendix E and G.
5.1 Impacts on Backbone Models
By introducing an additional 0.5% $\sim$ 1.3% (1.77M $\sim$ 33.5M) of parameters, HMT can enhance models with a variety of architectures to improve generation quality when processing long context inputs. We demonstrate this feature with general language modeling and question-answering tasks.
HMT consistently improves the backbone models in general language modeling tasks when processing long inputs. Figures 2 and 3 compare the perplexity of OPT 2.7B, RWKV 3B, and OpenLlamaV2 3B models with and without HMT on the Wikitext-103 and PG-19 datasets. Over input spanning from 2k $\sim$ 100k tokens, HMT consistently raises the generation quality of all these models. Moreover, Table 2 presents how improvements are achieved by HMT scales with the model size for same-family models. To further strengthen our argument that HMT can benefit larger models, we evaluate HMT with Qwen 2.5 14B utilizing 4 A100-80GB GPUs for training. As depicted in Figure 4, HMT can still increase the effectiveness of the backbone model on PG-19.
Notice that the improvement is not necessarily contributed solely by the additional parameters. Having more parameters does not always lead to higher performance. For example, HMT boosts OPT 2.7B to realize a lower perplexity than OpenLlama 3B with 20.7% fewer parameters, while OPT 2.7B performs worse without HMT. Section 5.2 describes more examples of HMT achieving superior generation quality with smaller models.
Table 2: Scalability of HMT. Average PPL is computed by taking the average PPL for samples in each sequence length in the experiment.
Model Avg Test PPL (Wikitext) ($\downarrow$)
OPT 350M 15.11
HMT + OPT 350M 14.28 (-5.8%)
OPT 2.7B 12.12
HMT + OPT 2.7B 8.61 (-28.9%)
RWKV 430M 19.33
HMT + RWKV 430M 16.10 (-16.6%)
RWKV 3B 13.30
HMT + RWKV 3B 9.93 (-25.3%)
HMT enhances long-answer contextual reasoning and short-answer prediction ability in question-answering tasks. One of the use cases of HMT is handling question-answering tasks that involve multiple contexts. Thus, we select the PubMedQA dataset and derive long-context QA samples with controllable context counts to evaluate the effectiveness of HMT. Two metrics are employed: for long answers, we compute the PPL to assess the contextual reasoning of HMT; for short answers, we measure the response accuracy. As seen in Figures 5 and 6, for samples with 2 to 10 contexts, HMT increases the effectiveness in PPL by 9.48% for long answers. For short answer tasks, HMT is 1.0% more accurate than the backbone model and exhibits significant advantages when samples have more contexts. In sum, HMT increases both the correctness and reasoning ability of models in long-context QA tasks.
Figure 4: Test Perplexity of HMT, RMT, and baseline model for Qwen 2.5 14B on PG-19 dataset. HMT boosts the effectiveness of the baseline model by 10.0%, while RMT worsens its effectiveness.
Figure 5: Long answer quality of RMT and HMT applied on Llama-2 7B, evaluated over PubMedQA dataset. HMT is 8.98% more effective than RMT.
Figure 6: Short response accuracy of RMT and HMT applied on Llama-2 7B, evaluated over PubMedQA dataset. HMT is 1.8% more accurate than RMT.
5.2 Comparison to Long Context Models
Combined with small and short-context models, HMT can be more effective than large models trained on long-context inputs. Table 3 displays metric results of HMT-augmented models on subsets of LongBench Bai et al. (2023b) and compares them with large models specialized for long contexts. The subsets contain various generation tasks, including single/multi-document QA, summarization, and few-shot learning. With a significantly lower inference memory requirement, HMT applied to small models can attain comparable or better metrics compared to large models, indicating a significant resource advantage. Specifically, we observe that HMT with small models performs well in generating short responses for long and multi-context inputs, thanks to its context-filtering ability. However, it exhibits comparable or weaker performance in generating long responses, as small models have shorter token generation limits compared to large models.
Moreover, applying HMT to long-context models can further improve their effectiveness and reduce inference memory consumption. For example, the AMD MI210 GPU cannot handle inferencing 30k token inputs with the Yi-6B-200K model due to memory constraints. Applying a sliding window strategy with a 5.2K-token window (Yi-6B-SW-5.2K), the model consumes 44.8 GB VRAM. On the contrary, HMT + Yi-6B-200K requires only 33.9 GB VRAM to process 30k tokens with a small segment length (512 tokens), with a 2% effectiveness improvement. Table 4 presents the effectiveness of long-range models on Wikitext-103 compared with several HMT-augmented models, including Mamba and Mistral models.
Table 3: Metric results of HMT-augmented small models and large models trained on longer contexts. Models with HMT can process infinitely long context, but only keep a fixed length of KV cache (the value in the parenthesis). We evaluate on subsets of LongBench, including QMSum (QMS), MuSiQue (MSQ), Qasper (QASP), NarrativeQA (NQA), MultiFieldQA-en (MFQA-en), GovReport (GR), TriviaQA (TQA), HotpotQA (HQA), 2WikiMQA (2WMQA), and MultiNews (MN). Mem Req indicates the minimum inference memory required (to store parameters and KV cache). Actual inference may require a larger VRAM.
Model Context Window QMS MSQ QASP NQA MFQA-en GR TQA HQA 2WMQA MN Mem Req (GB $\downarrow$)
GPT-3.5 Turbo 16384 23.4 26.9 43.3 23.6 52.3 29.5 91.4 51.6 37.7 26.7 46.7
Llama 2 7B Chat 4096 20.8 9.4 19.2 18.7 36.8 27.3 77.8 25.4 32.8 25.8 15.1
LongChat V1.5 7B 32768 22.7 9.7 27.7 16.9 41.4 30.8 82.3 31.5 20.6 26.4 22.6
XGen 7B 8192 20.5 10.3 18.1 18.0 37.7 27.3 77.8 29.7 21.1 26.2 16.2
InternLM 7B 8192 15.9 9.0 16.7 12.1 23.4 9.7 77.8 28.7 21.1 22.8 16.2
ChatGLM2 6B 32768 24.0 21.9 31.5 21.1 46.2 32.4 78.7 45.1 34.0 26.5 19.5
Vicuna V1.5 7B 16384 22.8 9.8 26.1 19.4 38.5 27.9 86.2 25.3 20.8 27.2 18.3
ChatGLM3 6B 32768 23.9 40.4 43.3 26.0 51.7 36.8 87.1 54.4 44.9 27.9 19.5
HMT + OPT 350M $\infty$ (1024) 22.2 15.6 32.3 17.2 62.7 13.8 87.2 25.3 29.2 10.5 0.75
HMT + OpenLlamaV2 3B $\infty$ (512) 21.4 42.3 35.7 21.6 58.1 15.8 93.0 29.9 27.7 11.0 6.1
HMT + SmolLM 135M $\infty$ (1024) 19.4 14.1 30.2 19.6 48.1 15.7 81.3 27.6 23.3 9.5 0.4
Benchmark Average Len (words) 10614 11214 3619 18409 4559 8734 8209 9151 4887 2113
Table 4: Quality of long context models and HMT with various backbone models. The input size is 30k tokens and the dataset is Wikitext-103.
Model Max Context Test PPL (Wikitext)
RWKV 3B $\infty$ 13.13
Mamba 370M $\infty$ 87.08
Yi-6B-200K 200K OOM  [^2] Although this model is trained with 200K-token samples, it cannot be run on MI210 due to memory constraints.
Yi-6B-SW-5.2K 200K 6.89
Mistral-7B 32K 5.47
HMT + OPT 350M $\infty$ (1024) 13.67
HMT + OpenLlamaV2 3B $\infty$ (512) 7.04
HMT + RWKV 3B $\infty$ (256) 10.94
HMT + Mamba 370M $\infty$ (256) 16.71
HMT + Yi-6B-200K $\infty$ (512) 6.75
HMT + Mistral-7B $\infty$ (512) 5.12
5.3 Comparison to Memory-augmented and Hierarchical Methods
One popular memory-augmented model is the recurrent memory transformer Bulatov et al. (2022) (RMT). Our assessment indicates that HMT is generally better at both language modeling and question-answering tasks than RMT, illustrated in Figures 2, 3, 5, and 6. The improvement gap is especially significant for recurrent models such as RWKV. HMT can further increase the effectiveness of RWKV while RMT will degrade the performance for both datasets, as demonstrated in Figure 3. Since RWKV has already compressed past tokens and passed hidden states along the sequence, applying RMT to RWKV re-weights past information compressed in states periodically. This was originally done by the time-mixing module of RWKV. Therefore, the advantage of memory augmentation is limited. Due to the gradient vanishing issue, the model is harder to train with RMT, leading to inferior performance. However, we believe that the memory retrieval mechanism in HMT helps RWKV to select previous hidden states with the most relevance, boosting its effectiveness. Another advantage of HMT over RMT is its scalability with large models: while RMT applied to Qwen 2.5 14B results in reduced effectiveness compared to direct inference with the backbone model, HMT continues to enhance effectiveness, as illustrated in Figure 4.
Furthermore, compared with other memory-augmented models, HMT is not only easy to use but also has higher generation quality. Table 5 picks three memory-augmented methods (Memorizing Transformer Wu et al. (2022), LongMem Wang et al. (2024), and CCM-concat Kim et al. (2023)) and compares them with HMT with the same or similar-sized backbone models. We choose the datasets used by the original works for fair comparisons. Memorizing transformer and LongMem require modifying the core architecture of the base model. Future models cannot easily adopt such modifications. Overall, HMT outperforms these methods. We also list the inference memory overhead complexity for each model, where $L$ is the total context length, $l_{i}$ is the inference segment length, $l_{m}$ is the memory size ($L>l_{m}>l_{i}$), and $t$ is the number of memory embeddings concatenated for CCM-concat. HMT has the lowest memory complexity over all previous methods.
Figure 7: Comparison between HMT and HOMER without context extension and with YaRN, all applying on Llama 2 7B. On average, HMT is 9.9% more effective than HOMER with YaRN on PG-19.
Lastly, we compare HMT with HOMER Song et al. (2024), a method that hierarchically compresses inputs to reduce their length for inference. In terms of memory complexity, HOMER requires $O(\log(L))$ memory to store the reduction tree, leading to increased peak memory utilization as input length grows. In contrast, HMT maintains a constant peak memory complexity regardless of input length. Regarding effectiveness, HMT achieves 9.9% lower perplexity on PG-19 compared to HOMER with YaRN Peng et al. (2023b) for context extension. As shown in Figure 7, the benefits of HMT become more substantial as input length increases, highlighting its superior scalability with longer inputs.
Table 5: Comparison between HMT with previous memory-augmented methods (Memorizing Transformer, LongMem, and CCM-concat).
Model Test PPL (Wikitext, 30k token) Mem Overhead
MemTRM 31.51 $O(L)$
HMT + OPT 350M 13.67 $O(l_{i})$
Model Test PPL (ArXiv, variable) Mem Overhead
LongMem 10.08 $O(l_{m})$
HMT + Qwen1.5-0.5B 9.02 $O(l_{i})$
Model Test PPL (PG-19, 60k token) Mem Overhead
CCM-concat 7.41 $O(t+l_{i})$
HMT + Llama 2 7B 7.40 $O(l_{i})$
6 Conclusion
We present HMT, a framework to augment models’ long-range language processing ability with context switching. Inspired by the brain’s memory hierarchy, HMT imitates human memorization behavior by deploying hierarchical memory and the memory retrieval mechanism. HMT consistently improves the generation quality of the backbone models. Compared with other long-context LLMs and memory-augmented models, HMT achieves higher generation quality with lower memory requirements. Our model provides LLM accessibility to resource-constrained applications and represents a step forward to lifelong language tasks.
7 Limitations and Ongoing Works
  • Currently, HMT will save $N$ memory embeddings for memory search, which is a cross-attention layer. When $N$ is small (e.g., $N=300$), which is already sufficient for 100k token samples, the overhead is negligible. However, when $N$ grows and the memory embeddings are stored in different physical memory hierarchies, the overhead can be significant. An intelligent memory prefetching mechanism can potentially alleviate the latency overhead, which we leave as future work.
  • Due to the large computational graph of models when training with BPTT, tuning the extra parameters introduced by HMT can be memory-consuming, impeding experiments on larger-scale models. A more efficient way to extend BPTT depth without memory overhead is a future research direction.
  • Although HMT employs only one level of long-term memory, one may use multiple levels of long-term memory to improve information access efficiency. Similar techniques have been used for multilevel optimization in VLSI physical design Cong and Shinnerl (2013); Chan et al. (2005).
8 Ethical Statements
The capability of memorizing information by HMT offers convenience to people’s daily lives, while also raising concerns about privacy leakage through conversation with language model agents. Nevertheless, with further efforts to deploy it on edge devices without network connections, this issue can be resolved.
Acknowledgement
Appendix A Other Related Works
The memory-augmented long-context transformer has been an active research topic in recent years. LongMem Wang et al. (2024) chunks the long-document input into segments and caches the attention keys and values for each segment. During the inference of a segment, LongMem will select relevant key-value embedding pairs by computing the attention score between token embeddings and the cached key embeddings and fuse the top k embeddings. Memorizing Transformer Wu et al. (2022) also caches the key-value embedding pairs similar to LongMem, but utilizes a kNN search to retrieve information similar to Unlimiformer. RET-LLM Modarressi et al. (2023) employs prompt engineering to store the informative context in a database and search keywords when the context involves questions. $\text{Memory}^{3}$ Yang et al. (2024) compresses segments of tokens into “explicit" memory blocks and stores them directly into a memory bank for retrieval. While these works can precisely retrieve contexts, they are not scalable due to the increasing memory consumption of storing long contexts without compression. Segment-level recurrent models, such as EMMA Moro et al. (2023), Memformer Wu et al. (2020), and Transformer-XL Dai et al. (2019), attempt to compress memory throughout the recurrence to reduce memory consumption. EMMA composes long-term memory from multiple short-term memory by linear combination and concatenates long and short-term memory to augment the segments. Transformer-XL propagates the compressed memory states derived from the attention of current layers to the previous layers for every iteration. Memformer augments the attention with the stored memory embeddings per time step and retrieves information using the cross-attention layer of the encoder-decoder model. However, Memformer employs a forgetting network to remove irrelevant context similar to LSTM, which can potentially delete useful contexts for unseen inputs. On the other hand, HMT condenses contexts into embeddings and retrieves information precisely without requiring a forgetting network to remove information permanently. Also, some of these works, including Memorizing Transformer, Memformer, TransformerXL, and Unlimiformer, need to fundamentally change the model architecture or inject new adapters based on different base model architecture. It makes deployment and extension to future LLMs very expensive. HMT avoids this issue by having a model-independent plug-and-play framework.
Appendix B Comparison to LongMem
Unlike HMT, LongMem Wang et al. (2024) operates on the key and value caches of each layer of the model and requires caching long caches to capture distant context. To compare with LongMem, we pick the Qwen1.5-0.5B Bai et al. (2023a) model as the backbone model and train HMT by 700 steps with 4 segments over 100 samples of the ArXiv subset of the Pile dataset Gao et al. (2020). The subset has 15.4K tokens on average and 60K tokens on maximum per sample, as Wang et al. (2024) described. Due to the large storage consumption of the training subset of the Pile dataset, we only extract the ArXiv subset in the validation and test split. HMT is trained on the validation set and tested on the test set. Table 6 illustrates that HMT + Qwen1.5-0.5B realizes lower PPL, with a smaller parameter size and in-memory length (number of memory embeddings, which is the number of key and value embeddings cached for LongMem). This indicates that HMT is memory efficient.
Table 6: Effectiveness of LongMem Wang et al. (2024) and HMT + Qwen1.5-0.5B models over ArXiv subset of the Pile dataset. With HMT, the Qwen1.5-0.5B model can obtain better effectiveness with fewer parameters and shorter memory, after 700 steps of update. The result for LongMem comes from the original paper. Subscription is the standard deviation.
Model # Params Segment Length In-memory Length Test PPL (ArXiv)
LongMem 558M 1k 65k 10.08
HMT + Qwen1.5-0.5B 463M 1k 300 $9.02_{\pm 0.04}$
Appendix C Comparison to Unlimiformer
There are two major differences between a previous retrieval-augmented model, Unlimiformer Bertsch et al. (2023), and HMT in terms of the memory retrieval mechanism:
  • Unlimiformer retrieves the information with kNN search over the collection of encoded token segments, while HMT uses cross-attention. We believe there are several advantages of employing cross-attention: (1) Attending Top K’s most similar token segments still introduces information loss. Regarding the self-attention layer, the aggregation of tokens with less similar encodings may positively contribute to the quality of the final output. On the other hand, cross-attention fuses all cached hidden embeddings, weighted by the relative similarity, which captures the whole context. (2) The output of the cross-attention is a single embedding, which has lower computational overhead compared to attending k extra tokens.
In terms of usage, Unlimiformer targets encoder-decoder models and injects retrieval modules into the backbone model. Although the authors recently added support for decoder-only models for token generation, only the Llama model architectures Touvron et al. (2023) can be applied and the training/evaluation procedure is not specified. This is one of the biggest challenges for Unlimiformer to adapt to future LLMs for validation and generation. On the contrary, HMT focuses on decoder-only models. Since HMT does not inject new modules in the backbone model, it is effortless to adapt to future LLMs.
Appendix D HMT, RMT, and Baseline Training Details and Hyperparameters
Table 7 are the training configurations of the backbone models and HMT/RMT.